# A Paradigm for Creative Ownership

### v1

@@ 4. Future Work
The next steps for this work involve conducting a user study with creative practitioners and professionals to understand how closely the proposed framework aligns with the experiences of creatives in a range of fields. To do so, we will ask practitioners how strongly they feel ownership sentiments over a creative artifact that they have produced before walking them through the questions on the framework. We hope to understand whether reported feelings of ownership align with the results of the framework, whether some elements are more important than others or if there are any missing elements, and whether there are differences between various fields of practice. Following validation of the framework, we envision this as being a useful research tool for studying ownership in human-AI collaboration contexts.
Acknowledgements.
Heartfelt gratitude to Ami Mehta, Samuel Budnyk, Victor Arul, and Flora Saki Giordani for generously sharing their expertise and insights as artists and creators. Their contributions have been invaluable to this work.

### v2

@@ 9. Future Work
Rather than treating dimensions independently, future work should examine their interactions. For example, we repeatedly observed Control entangled with both Intentionality (ability to execute one’s vision) and Embodiment (expressing identity/values). A second direction is within-domain depth: interviewing a cohort within a single field (e.g., 20 choreographers, or 20 ceramicists) can surface nuances and interaction patterns obscured by cross-field sampling. Finally, education contexts invite domain-specific hypotheses: in creative writing, Embodiment may be primary, whereas in mathematics instructors may prioritize Intentionality. The multidimensional nature of this framework sheds light on these differences when it comes to how we think about evaluating the impacts of a system.
10. Conclusion
We introduced a nine-subdimension framework for creative ownership and an interactive web tool, and evaluated both through interviews with 21 creative professionals. The framework broadened what participants discussed beyond a few familiar themes and separated high- from low-ownership cases while revealing heterogeneity in how ownership diminishes.
For HCI, the immediate use is practical: report ownership as a profile rather than a single score, state construct boundaries, and use the dimensions as design levers (e.g., decision rights for Control, intent alignment for Intentionality, attribution for Recognition, modality-aware workflows for Production/Abstraction, and role clarity for Interdependence).
Next steps include testing interactions among dimensions, running within-domain and longitudinal studies, and using the framework as an outcome measure when evaluating AI-in-the-loop tools. Adoption of the framework and tool should make ownership findings more comparable across studies and help designers build systems that support, rather than erode, felt ownership.
@@ Appendix C Creative Ownership Framework Applications
Appendix D Interview Guide
This appendix reproduces the interview protocol approved by our Institutional Review Board (IRB). Participants were recruited through professional contacts, publicly available sources, and snowball sampling, and were screened to ensure that they: (1) work or participate significantly in a creative field, (2) have at least two finished creative products they are willing to discuss, (3) are fluent in English, and (4) are over 18 years of age.
Interviews lasted approximately 45–60 minutes. At the beginning of each interview, the researcher reiterated the study goals, confirmed that the participant had read the consent form, and obtained verbal consent to record the session. Interviews followed a semi-structured protocol with the following guiding questions and activities:
  1. (1)
    Can you tell us about your creative background, current field of work, and expertise?
  2. (2)
    Can you tell us about two creative products—one associated with low feelings of ownership and one associated with high feelings of ownership—that you have worked on?
  3. (3)
    Please rank your overall feelings of ownership over each of these projects on a scale of 1–10.
  4. (4)
    How did you arrive at these ratings? What aspects of the project evoke these feelings?
  5. (5)
    At this step, participants will be asked to fill out the Creative Ownership Test survey for each of the creative products they have worked on.
  6. (6)
    Do you feel that the results of this test align with your original ownership feelings?
  7. (7)
    Are there any dimensions of this framework that feel particularly relevant to your feelings of ownership?
  8. (8)
    Are there any dimensions that are uninformative, redundant, or not useful toward your feelings of ownership?
  9. (9)
    Are there any elements you feel are missing from this framework that could better explain your ownership feelings?
Interviews followed a semi-structured format in which participants answered guiding questions and, where relevant, engaged in follow-up discussion to elaborate on their responses. Data were analyzed anonymously, with only a participant number and their creative practice used as identifiers.